OpenStudy (anonymous):

Solve by the substitution method 5x-6y=-12 5x+52=y

OpenStudy (anonymous):

\[5x-6(5x+52)=-12\]

OpenStudy (anonymous):

First you will need to rewrite the equation so that one variable is alone (no coefficient) and is on one side by itself. The second equation has y already by itself. Now we just plug 5x-52 in for Y in the first equation to get what Lammy has posted

OpenStudy (anonymous):

So confused

OpenStudy (anonymous):

I can draw this out for you (literally)

OpenStudy (anonymous):

\[5x-30X-312=-12 \]

OpenStudy (anonymous):

Go here and I can explain in detail http://flockdraw.com/8v51mc

OpenStudy (anonymous):

you have y=5x + 52 just plug that into y on the top equation

OpenStudy (anonymous):

and then use distributive properties

OpenStudy (anonymous):

hey sorry for drawing on the thing. i was trying to chat. but couldnt figure out how.

OpenStudy (anonymous):

its a cool software though.

OpenStudy (anonymous):

it is you just click the A

OpenStudy (anonymous):

This is actually even easier than that.. \[5x-6y=-12\]\[\implies 5x = 6y-12\] \[5x+52=y\]\[\implies (6y-12) + 52 = y\] No distribution needed.

OpenStudy (anonymous):

thanks pol pak

OpenStudy (anonymous):

Not that you should avoid learning to distribute well. =)

OpenStudy (anonymous):

thank you I am learning as we speak. but thank you!!! I am trying to learn algebra...I am lol